# Smooth orbit equivalence rigidity for dissipative geodesic flows

### v1

@@ 4.1. Torelli’s theorem
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be Anosov thermostats satisfying the attenuated tensor tomography problem of order $1$. If there exists a smooth orbit equivalence isotopic to the identity between them, then $[J]=[\tilde{J}]$ in $\mathcal{M}(M)$. Equivalently, there exists a diffeomorphism $\psi:M\to M$ such that $\psi^{*}\tilde{J}=J$ and $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$.
@@
By Proposition 3.4, the map $\Psi:H^{0}_{J}(M,\kappa)\to H^{0}_{\tilde{J}}(M,\kappa)$ is a period-preserving $\mathbb{C}$-linear isomorphism. This means that $(M,J)$ and $(M,\tilde{J})$ have the same period matrix. Indeed, given a canonical basis $\{a_{j},b_{j}\}$ of the homology $H_{1}(M;\mathbb{Z})$ on $M$, let $\{\zeta_{j}\}$ be a basis for $H_{J}^{0}(M,\kappa)$ such that property (2.7) is satisfied. Then $\{F\zeta_{j}\}$ is a basis for $H^{0}_{\tilde{J}}(M,\kappa)$ such that (2.7) is also satisfied, and
$(\pi(J))_{jk}=\int_{b_{j}}\zeta_{k}=\int_{b_{j}}F\zeta_{k}=(\pi(\tilde{J}))_{jk}.$
Since the surface $M$ must be of genus $\geq 2$, Theorem 2.8 tells us that there exists an orientation-preserving diffeomorphism $\psi:M\to M$ such that $\psi^{*}J=\tilde{J}$.
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ be either
@@
    two Anosov magnetic systems, or
@@
    two Gaussian thermostats with $\mathbb{K},\tilde{\mathbb{K}}<0$.
If there exists a smooth orbit equivalence isotopic to the identity between them, then $[J]=[\tilde{J}]$ in $\mathcal{T}(M)$. Equivalently, there exists a diffeomorphism $\psi:M\to M$, isotopic to the identity, such that $\psi^{*}\tilde{J}=J$ and $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$.
The reason we needed to specify the nature of the two thermostats, as opposed to Proposition 4.1 which deals with more general Anosov thermostats of Fourier degree $\leq 1$, is that our proof relies the following technical lemma (see [GLP24, Lemma 3.8]).
@@
Let $J$ and $\tilde{J}$ be two complex structures on $M$ compatible with orientation such that $[J]\neq[\tilde{J}]$ in $\mathcal{T}(M)$. Then, there exists a finite cover $M^{\prime}\to M$ such that the lifts $[J^{\prime}],[\tilde{J}^{\prime}]\in\mathcal{T}(M^{\prime})$ are not in the same $\textup{MCG}(M^{\prime})$-orbit.
Indeed, when lifting the thermostats to finite covers, the properties of being Anosov and having negative thermostat curvature are preserved, but, a priori, satisfying the attenuated tensor tomography problem of order $1$ is not.
@@
Suppose, for the sake of contradiction, that $[J]\neq[\tilde{J}]$ in $\mathcal{T}(M)$. By Lemma 4.3, there exists a finite cover $M^{\prime}\to M$ such that the lifts $[J^{\prime}]$ and $[\tilde{J}^{\prime}]$ are not in the same $\text{MCG}(M^{\prime})$-orbit.
@@
Proposition 4.2 gives us most of Theorem 1.1 and Theorem 1.2. All that is left is studying the rigidity of the function $\lambda$ in each case.
@@ 4.2. Rigidity of the magnetic field
Since $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ does not depend on the velocity in the magnetic case, we will think of it as living in $\mathcal{C}^{\infty}(M,\mathbb{R})$.
@@
Let $(M,g,\lambda)$ be an Anosov magnetic flow. Then, the $2$-form $\omega$ on $SM$ defined in (2.2) is exact.
@@
By the Gauss-Bonnet theorem and the fact $M$ must be of genus $\geq 2$, we know that
@@
so $[K_{g}\mu_{a}]\neq 0$ in $H^{2}(M;\mathbb{R})\cong\mathbb{R}$. It follows that we may write $\lambda\mu_{a}=cK_{g}\mu_{a}+d\varrho$ for some $1$-form $\varrho$ on $M$ and $c\in\mathbb{R}$. The constant $c$ is explicitly given by
@@
But then, since $d\psi=-\pi^{*}(K_{g}\mu_{a})$, we obtain
@@
$\tau:=-\alpha-c\psi+\pi^{*}\varrho.$ (4.2)
Knowing how to find primitives of $\omega$ in the magnetic case then unlocks the following.
@@
Let $(M,g,\lambda)$ and $(M,g,\tilde{\lambda})$ be two Anosov magnetic systems with the same background metric $g$. Suppose there is a smooth conjugacy $\phi:SM\to SM$, isotopic to the identity, between them. Then, we have $[\tilde{\lambda}\mu_{a}]=\pm[\lambda\mu_{a}]$ in $H^{2}(M;\mathbb{R})$. Moreover, $\lambda=0$ if and only if $\tilde{\lambda}=0$.
@@
Define the $\tau$ as in (4.2) to be a primitive of $\omega$. Contracting it with $F$ yields
$\tau(F)=-1-c\pi^{*}\lambda+\pi^{*}_{1}\varrho.$ (4.3)
Moreover, we know that the Anosov magnetic flows are transitive and $\phi$ is isotopic to the identity, so $\phi^{*}\mu=\mu$. Since $\phi_{*}\tilde{F}=F$, contracting $\phi^{*}\mu=\mu$ with $\tilde{F}$ yields $\phi^{*}\omega=\tilde{\omega}$. This can be rewritten as $\phi^{*}d\tau=d\tilde{\tau}$, which in turn implies
@@
Since $\pi^{*}:H^{1}(M;\mathbb{R})\to H^{1}(SM;\mathbb{R})$ is an isomorphism, there exists a closed $1$-form $\varphi$ on $M$ and $f\in\mathcal{C}^{\infty}(SM)$ such that
@@
Contracting with $\tilde{F}$ yields
$\tau(F)\circ\phi=\tilde{\tau}(\tilde{F})+\pi_{1}^{*}\varphi+\tilde{F}f.$
By (4.3), we thus get
$-1+(-c\pi^{*}\lambda+\pi_{1}^{*}\varrho)\circ\phi=-1-\tilde{c}\pi^{*}\tilde{\lambda}+\pi_{1}^{*}\tilde{\varrho}+\pi_{1}^{*}\varphi+\tilde{F}f,$
@@
$(-c\pi^{*}\lambda+\pi_{1}^{*}\varrho)\circ\phi=-\tilde{c}\pi^{*}\tilde{\lambda}+\pi_{1}^{*}\tilde{\varrho}+\pi_{1}^{*}\varphi+\tilde{F}f.$ (4.4)
@@
We thus have $c^{2}=\tilde{c}^{2}$ by (4.1), which shows that the cohomology classes of $\lambda\mu_{a}$ and $\tilde{\lambda}\mu_{a}$ in $H^{2}(M;\mathbb{R})$ are the same up to a sign.
If $\lambda=0$, we may take $\varrho=0$, and we know that $c=0$ thanks to (4.1). It follows that $\tilde{c}=0$. Let $\tilde{\gamma}$ be a closed magnetic geodesic of $(M,g,\tilde{\lambda})$ and $\tilde{\Gamma}:=(\tilde{\gamma},\dot{\tilde{\gamma}})\subseteq SM$. Relation (4.4) allows us to write
$\int_{\tilde{\Gamma}}\pi_{1}^{*}(\tilde{\varrho}+\varphi)=0.$
By the smooth Livšic theorem [dlLMM86, Theorem 2.1] and [DP05, Theorem B], this means that $\tilde{\varrho}+\varphi$ is exact. Since $\varphi$ is closed, we get $d\tilde{\varrho}=0$, which in turn implies $\tilde{\lambda}=0$, as desired. ∎
@@
If two Anosov magnetic systems $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ are related by a smooth conjugacy $\phi:\tilde{S}M\to SM$, isotopic to the identity, Proposition 4.2 yields a diffeomorphism $\psi:M\to M$ isotopic to the identity such that $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$.
If $\phi$ is a conjugacy and $f=0$, the map $\phi\circ d\psi:SM\to SM$ gives us a smooth conjugacy isotopic to the identity between the Anosov magnetic flows $(M,g,\psi^{*}\tilde{\lambda})$ and $(M,g,\lambda)$. Thus Proposition 4.5 tells us that $[\psi^{*}(\tilde{\lambda})\mu_{a}]=\pm[\lambda\mu_{a}]$ in $H^{2}(M;\mathbb{R})$ and that $\lambda=0$ if and only if $\tilde{\lambda}=0$.
@@ 4.3. Rigidity of the thermostat 1-form
Given the conclusion of Proposition 4.2, it behooves us to understand the behavior of (generalized) thermostat flows under a conformal re-scaling of the metric.
@@
Let $(M,g,\lambda)$ be a generalized thermostat, and define a conformal re-scaling $\tilde{g}:=e^{-2f}g$ of the metric, for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$. Then, the scaling map $s:SM\to\tilde{S}M$ defined in (1.1), which in this case is simply $(x,v)\mapsto(x,e^{f}v)$, satisfies
$s_{*}X=e^{-f}(\tilde{X}-\tilde{\pi}_{1}^{*}(\star df)\tilde{V}),\qquad s_{*}V=\tilde{V}.$
In particular, the map $s$ represents a smooth orbit equivalence isotopic to the identity from $(M,g,\lambda)$ to
$(M,\tilde{g},e^{f}(\lambda\circ s^{-1})-\tilde{\pi}_{1}^{*}(\star df))$ (4.5)
with a time-change $s_{*}F=e^{-f}\tilde{F}$.
@@
The first statement is proved in [CP22, Lemma B.1]. The conclusion then follows from the calculation
@@
$=e^{-f}(\tilde{X}-\tilde{\pi}_{1}^{*}(\star df)\tilde{V})+(\lambda\circ s^{-1})\tilde{V}.$
In what follows, let $\theta$ be the $1$-form on $M$ satisfying $\pi_{1}^{*}\theta=\lambda_{-1}+\lambda_{1}$. If $\lambda$ is of Fourier degree $\leq 1$, we may more succinctly write the thermostat (4.5) as
$(M,e^{-2f}g,e^{f}\lambda_{0}+\tilde{\pi}_{1}^{*}(\theta-\star df)).$
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be two Anosov thermostats. Suppose there is a smooth orbit equivalence $\phi:\tilde{S}M\to SM$, isotopic to the identity, between them. If $\star\theta$ or ~$\tilde{\star}\tilde{\theta}$ is closed, then $\star\theta-\tilde{\star}\tilde{\theta}$ is exact.
@@
By Lemma 2.1, we have $\mathrm{div}_{\mu}F=V(\lambda)=-\pi_{1}^{*}(\star\theta)$, so an application of Lemma 2.2 gives us
$(\phi^{*}c)\phi^{*}(\pi_{1}^{*}(\star\theta))=\tilde{\pi}_{1}^{*}(\tilde{\star}\tilde{\theta})-\tilde{F}\left(\ln\left(\dfrac{\det\phi}{\phi^{*}c}\right)\right),$
where $c\in\mathcal{C}^{\infty}(SM,\mathbb{R}_{>0})$ is such that $\phi_{*}\tilde{F}=cF$. Therefore, if $\tilde{\gamma}$ is a closed thermostat geodesic on $M$ with respect to $(M,\tilde{g},\tilde{\theta})$, and $\tilde{\Gamma}:=(\tilde{\gamma},\dot{\tilde{\gamma}})\subseteq\tilde{S}M$, we have
$\int_{\phi(\tilde{\Gamma})}c\pi^{*}_{1}(\star\theta)=\int_{\tilde{\Gamma}}\phi^{*}(c\pi^{*}_{1}(\star\theta))=\int_{\tilde{\Gamma}}\tilde{\pi}^{*}_{1}(\tilde{\star}\tilde{\theta}).$
@@
$\int_{\phi(\tilde{\Gamma})}c\pi^{*}_{1}(\star\theta)=\int_{\tilde{\Gamma}}\tilde{\pi}_{1}^{*}(\star\theta).$
@@
$\int_{\tilde{\Gamma}}\tilde{\pi}_{1}^{*}(\star\theta-\tilde{\star}\tilde{\theta})=0.$
An application of the smooth Livšic theorem [dlLMM86, Theorem 2.1] and [DP07, Theorem B] allows us to conclude that $\star\theta-\tilde{\star}\tilde{\theta}$ is exact, as desired.
@@
It remains to show that, if either $\star\theta$ or ~$\tilde{\star}\tilde{\theta}$ is closed, then $\star(\psi^{*}\tilde{\theta}-\theta)$ is exact. The map $\phi\circ d\psi$ yields a smooth orbit equivalence isotopic to the identity between the Anosov Gaussian thermostats $(M,e^{2f}g,\psi^{*}\tilde{\theta})$ and $(M,g,\theta)$. By Lemma 4.6, we may assume without loss of generality that $f=0$. Proposition 4.7 then allows us to conclude.

### v2

@@ 4.1. Torelli’s theorem
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be Anosov thermostats satisfying the attenuated tensor tomography problem of order $1$. If there exists a smooth orbit equivalence isotopic to the identity between them, then $[J]=[\widetilde{J}]$ in $\mathcal{M}(M)$. Equivalently, there exists a diffeomorphism $\psi:M\to M$ such that $\psi^{*}\widetilde{J}=J$ and $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$.
@@
By Proposition 3.4, the map $\Psi:H^{0}_{J}(M,\kappa)\to H^{0}_{\widetilde{J}}(M,\kappa)$ is a period-preserving $\mathbb{C}$-linear isomorphism. This means that $(M,J)$ and $(M,\widetilde{J})$ have the same period matrix. Indeed, given a canonical basis $\{a_{j},b_{j}\}$ of the homology $H_{1}(M;\mathbb{Z})$ on $M$, let $\{\zeta_{j}\}$ be a basis for $H_{J}^{0}(M,\kappa)$ such that property (2.7) is satisfied. Then, $\{\Psi\zeta_{j}\}$ is a basis for $H^{0}_{\widetilde{J}}(M,\kappa)$ such that property (2.7) is also satisfied, and
$(\Pi(J))_{jk}=\int_{b_{j}}\zeta_{k}=\int_{b_{j}}\Psi\zeta_{k}=(\Pi(\widetilde{J}))_{jk}.$
Since the surface $M$ must be of genus $\geq 2$, Theorem 2.8 tells us that there exists an orientation-preserving diffeomorphism $\psi:M\to M$ such that $\psi^{*}J=\widetilde{J}$. ∎
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ be either:
@@
    two Anosov magnetic systems; or
@@
    two Gaussian thermostats with $\mathbb{K},\widetilde{\mathbb{K}}<0$.
If there exists a smooth orbit equivalence isotopic to the identity between them, then $[J]=[\widetilde{J}]$ in $\mathcal{T}(M)$. Equivalently, there exists a diffeomorphism $\psi:M\to M$ isotopic to the identity such that $\psi^{*}\widetilde{J}=J$ and $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$.
The reason we need to specify the nature of the two thermostats, as opposed to Proposition 4.1 which deals with general Anosov thermostats of Fourier degree $\leq 1$, is that our proof relies on the following technical lemma (see [GLP25, Lemma 3.8]).
@@
Let $J$ and $\widetilde{J}$ be two complex structures on $M$ compatible with orientation such that $[J]\neq[\widetilde{J}]$ in $\mathcal{T}(M)$. Then, there exists a finite cover $M^{\prime}\to M$ such that the lifts $[J^{\prime}],[\widetilde{J}^{\prime}]\in\mathcal{T}(M^{\prime})$ are not in the same $\textup{MCG}(M^{\prime})$-orbit.
Indeed, when lifting thermostats to finite covers, the Anosov property and negative thermostat curvature are preserved; however, satisfying the attenuated tensor tomography problem of order $1$ is not preserved a priori.
@@
Suppose, for the sake of contradiction, that $[J]\neq[\widetilde{J}]$ in $\mathcal{T}(M)$. By Lemma 4.3, there exists a finite cover $p:M^{\prime}\to M$ such that the lifts $[J^{\prime}]$ and $[\widetilde{J}^{\prime}]$ are not in the same $\text{MCG}(M^{\prime})$-orbit.
@@
Proposition 4.2 gives us most of Theorems 1.2 and 1.3. All that is left is studying the rigidity of the function $\lambda$ in each case.
@@ 4.2. Rigidity of the magnetic field
Since $\lambda\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ does not depend on velocity in the magnetic case, we will think of it as living in $\mathcal{C}^{\infty}(M,\mathbb{R})$.
@@
Let $(M,g,\lambda)$ be an Anosov magnetic system. Then, the $2$-form $\omega$ on $SM$ defined in equation (2.2) is exact.
@@
By the Gauss–Bonnet theorem and the fact $M$ must be of genus $\geq 2$, we know that
@@
so $[K_{g}\mu_{a}]\neq 0$ in $H^{2}(M;\mathbb{R})\cong\mathbb{R}$. It follows that we may write $\lambda\mu_{a}=cK_{g}\mu_{a}+d\varrho$ for some $1$-form $\varrho$ on $M$ and constant $c\in\mathbb{R}$. The constant $c$ is explicitly given by
@@
However then, since $d\psi=-\pi^{*}(K_{g}\mu_{a})$, we obtain
@@
$\tau\coloneqq-\alpha-c\psi+\pi^{*}\varrho,$ (4.2)
as desired. ∎
Knowing how to find primitives of $\omega$ in the magnetic case then unlocks the following (compare to [Pat06, Lemma 4.1]).
@@
Let $(M,g,\lambda)$ and $(M,g,\tilde{\lambda})$ be two Anosov magnetic systems with the same background metric $g$. Suppose that there is a smooth conjugacy $\phi:SM\to SM$ isotopic to the identity between them. Then, we have $[\tilde{\lambda}\mu_{a}]=\pm[\lambda\mu_{a}]$ in $H^{2}(M;\mathbb{R})$. Moreover, $\lambda=0$ if and only if $\tilde{\lambda}=0$.
@@
Define $\tau$ as in equation (4.2) to be a primitive of $\omega$. Contracting it with $F$ yields
$\tau(F)=-1-c\pi^{*}\lambda+\ell_{1}\varrho.$ (4.3)
Recall that the restriction map $\ell_{1}$ is defined in equation (1.2). Moreover, we know that the Anosov magnetic flows are transitive and $\phi$ is isotopic to the identity, so $\phi^{*}\mu=\mu$. Since $\phi_{*}\widetilde{F}=F$, contracting $\phi^{*}\mu=\mu$ with $\widetilde{F}$ yields $\phi^{*}\omega=\tilde{\omega}$. This can be rewritten as $\phi^{*}(d\tau)=d\tilde{\tau}$, which in turn implies
@@
Since $\pi^{*}:H^{1}(M;\mathbb{R})\to H^{1}(SM;\mathbb{R})$ is an isomorphism, there exist a closed $1$-form $\varphi$ on $M$ and a function $f\in\mathcal{C}^{\infty}(SM,\mathbb{R})$ such that
@@
Contracting with $\widetilde{F}$ yields
$\tau(F)\circ\phi=\tilde{\tau}(\widetilde{F})+\ell_{1}\varphi+\widetilde{F}f.$
By equation (4.3), we thus get
$-1+(-c\pi^{*}\lambda+\ell_{1}\varrho)\circ\phi=-1-\tilde{c}\pi^{*}\tilde{\lambda}+\ell_{1}\tilde{\varrho}+\ell_{1}\varphi+\widetilde{F}f,$
@@
$(-c\pi^{*}\lambda+\ell_{1}\varrho)\circ\phi=-\tilde{c}\pi^{*}\tilde{\lambda}+\ell_{1}\tilde{\varrho}+\ell_{1}\varphi+\widetilde{F}f.$ (4.4)
@@
We thus have $c^{2}=\tilde{c}^{2}$ by equation (4.1), which shows that the cohomology classes of $\lambda\mu_{a}$ and $\tilde{\lambda}\mu_{a}$ in $H^{2}(M;\mathbb{R})$ are the same up to a sign.
If $\lambda=0$, we may take $\varrho=0$, and we know that $c=0$ thanks to equation (4.1). It follows that $\tilde{c}=0$. Let $\tilde{\gamma}$ be a closed magnetic geodesic for $(M,g,\tilde{\lambda})$ of period $T$ and let $\widetilde{\Gamma}\coloneqq(\tilde{\gamma},\dot{\tilde{\gamma}})\subset SM$. Equation (4.4) allows us to write
[…]
By the smooth Livšic theorem [dlLMM86, Theorem 2.1] and [DP05, Theorem B], this means that $\tilde{\varrho}+\varphi$ is exact. Since $\varphi$ is closed, we get $d\tilde{\varrho}=0$, which in turn implies that $\tilde{\lambda}=0$, as desired. ∎
@@
If two Anosov magnetic systems $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$ are related by a smooth conjugacy $\phi:\widetilde{S}M\to SM$ isotopic to the identity, Proposition 4.2 yields a diffeomorphism $\psi:M\to M$ isotopic to the identity such that $\psi^{*}\tilde{g}=e^{2f}g$ for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$.
If $\phi$ is a conjugacy and $f=0$, the map $\phi\circ d\psi:SM\to SM$ gives us a smooth conjugacy isotopic to the identity between the Anosov magnetic flows $(M,g,\psi^{*}\tilde{\lambda})$ and $(M,g,\lambda)$. Thus, Proposition 4.5 tells us that $[\psi^{*}(\tilde{\lambda})\mu_{a}]=\pm[\lambda\mu_{a}]$ in $H^{2}(M;\mathbb{R})$ and that $\lambda=0$ if and only if $\tilde{\lambda}=0$. ∎
@@ 4.3. Rigidity of the thermostat 1-form
Given the conclusion of Proposition 4.2, it behooves us to understand the behaviour of thermostat flows under a conformal rescaling of the metric.
@@
Let $(M,g,\lambda)$ be a thermostat and define a conformal rescaling $\tilde{g}\coloneqq e^{-2f}g$ of the metric for some $f\in\mathcal{C}^{\infty}(M,\mathbb{R})$. Then, the scaling map $s:SM\to\widetilde{S}M$ defined in equation (1.1), which in this case is simply $(x,v)\mapsto(x,e^{f}v)$, satisfies
$s_{*}X=e^{-f}(\widetilde{X}-\tilde{\ell}_{1}(\star df)\widetilde{V}),\qquad s_{*}V=\widetilde{V}.$
In particular, the map $s$ represents a smooth orbit equivalence isotopic to the identity from the thermostat $(M,g,\lambda)$ to the thermostat
$(M,\tilde{g},e^{f}(\lambda\circ s^{-1})-\tilde{\ell}_{1}(\star df)),$ (4.5)
with a time-change $s_{*}F=e^{-f}\widetilde{F}$.
@@
The first statement is proved in [CP25, Lemma B.1]. The conclusion then follows from the calculation
@@
$=e^{-f}(\widetilde{X}-\tilde{\ell}_{1}(\star df)\widetilde{V})+(\lambda\circ s^{-1})\widetilde{V}.\qed$
In what follows, let $\theta$ be the $1$-form on $M$ satisfying $\ell_{1}\theta=\lambda_{-1}+\lambda_{1}$. If $\lambda$ is of Fourier degree $\leq 1$, we may more succinctly write the thermostat (4.5) as
$(M,e^{-2f}g,e^{f}\lambda_{0}+\tilde{\ell}_{1}(\theta-\star df)).$
@@
Let $(M,g,\lambda)$ and $(M,\tilde{g},\tilde{\lambda})$, with $\lambda$ and $\tilde{\lambda}$ of Fourier degree $\leq 1$, be two Anosov thermostats. Suppose there is a smooth orbit equivalence $\phi:\widetilde{S}M\to SM$ isotopic to the identity between them. If $\star\theta$ or ~$\tilde{\star}\tilde{\theta}$ is closed, then $\star\theta-\tilde{\star}\tilde{\theta}$ is exact.
@@
By Lemma 2.1, we have $\mathrm{div}_{\mu}F=V\lambda=-\ell_{1}(\star\theta)$, so an application of Lemma 2.2 gives us
$\phi^{*}(c)\phi^{*}(\ell_{1}(\star\theta))=\tilde{\ell}_{1}(\tilde{\star}\tilde{\theta})-\widetilde{F}\left(\ln\left(\dfrac{\det\phi}{\phi^{*}c}\right)\right),$
where $c\in\mathcal{C}^{\infty}(SM,\mathbb{R}_{>0})$ is such that $\phi_{*}\widetilde{F}=cF$. Therefore, if $\tilde{\gamma}$ is a closed thermostat geodesic for $(M,\tilde{g},\tilde{\theta})$ o period $T$ and $\widetilde{\Gamma}\coloneqq(\tilde{\gamma},\dot{\tilde{\gamma}})\subset\widetilde{S}M$, we have
$\int_{0}^{T}\phi^{*}(c\ell_{1}(\star\theta))(\widetilde{\Gamma}(t))\,dt=\int_{0}^{T}\tilde{\ell}_{1}(\tilde{\star}\tilde{\theta})(\widetilde{\Gamma}(t))\,dt.$
@@
$\int_{0}^{T}\phi^{*}(c\ell_{1}(\star\theta))(\widetilde{\Gamma}(t))\,dt=\int_{0}^{T}\tilde{\ell}_{1}(\star\theta)(\widetilde{\Gamma}(t))\,dt.$
@@
$\int_{0}^{T}\tilde{\ell}_{1}(\star\theta-\tilde{\star}\tilde{\theta})(\widetilde{\Gamma}(t))\,dt=0.$
An application of the smooth Livšic theorem [dlLMM86, Theorem 2.1] and [DP07, Theorem B] allows us to conclude that $\star\theta-\tilde{\star}\tilde{\theta}$ is exact, as desired. ∎
@@
It remains to show that, if either $\star\theta$ or ~$\tilde{\star}\tilde{\theta}$ is closed, then $\star(\psi^{*}\tilde{\theta}-\theta)$ is exact. The map $\phi\circ d\psi$ yields a smooth orbit equivalence isotopic to the identity between the Anosov Gaussian thermostats $(M,e^{2f}g,\psi^{*}\tilde{\theta})$ and $(M,g,\theta)$. By Lemma 4.6, we may assume without loss of generality that $f=0$. Proposition 4.7 then allows us to conclude. ∎